# Distance-Domain Degrees of Freedom in Near-Field Region

### v3

@@ IV-B Eigenvalue Analysis via Fourier Transform Analysis
Scaled Fourier transform defined in (33) of $g(\Delta t)$ in (31) is given by:
@@
Proof: Proof is provided in Appendix B.
Figure 3: Spectrum of $\tilde{g}(\xi)$ for a squared Tx array $\mathcal{P}$ with a circular gap (radius = $60\lambda$) (Rx array spans over $r_{\min},r_{\max}]=[200\lambda,2000\lambda]$, and Tx array is bounded by ${p}_{\min},{p}_{\max}]=[60\lambda,100\lambda]$).
@@
$\text{DoF}\approx\frac{\left(p_{\max}^{2}-p_{\min}^{2}\right)}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1),$ (38)
where $O(1)$ denotes the negligible contribution from the sidelobes of the sinc function.
@@ IV-C1 Comparison with Angular-Domain DoF Expressions
In many P2P MIMO setups where the Rx aperture mainly spans the angular extent of the Tx aperture, the angular-domain DoF expressions often take the form [torkildson2011indoor, pizzo2022landau, miller2000communicating, xie2023performance, pu2014effects, ruiz2023degrees, do2023parabolic, xu2017degrees, yuan2021electromagnetic, xu2023exploiting, guo2025impact]:
@@
where $A_{t}$ and $A_{r}$ are the physical areas of Tx and Rx arrays, $D$ is the Tx–Rx separation (treated as a single representative distance), $n$ depends on aperture dimensionality, and $\alpha$ is coefficients determined via orientation of Tx and Rx arrays. More elaborate variants appear in [wang2025analytical]. Our distance-domain result in (38) differs in several key ways:
Figure 4: Normalized eigenvalues for three different shapes of Tx array with the same ${p}_{\max}=100\lambda,{p}_{\min}=60\lambda$.
1a. Effect of the Tx aperture size:
  • Angular-domain DoF (literature): DoF scales with aperture area (e.g., linearly with ULA length, quadratically with URA side length), since angular coverage grows with physical extent.
  • Distance-domain DoF (this work): DoF depends on $\big(p_{\max}^{2}-p_{\min}^{2}\big)$, i.e., only on the radial extremes of $\mathcal{P}$, regardless of its detailed shape:
    $\mathrm{DoF}\;\propto\;p_{\max}^{2}-p_{\min}^{2}.$
    For example, arrays with very different shapes (and areas) but the same $(p_{\max},p_{\min})$ yield nearly identical distance-domain DoF, as shown in Fig. 4. Consequently, for fixed area, elongating the aperture to increase $p_{\max}$ can be more effective than uniformly enlarging area (see Fig. 5 where we simulate three setups with the same area, i.e., $50\lambda\times 30\lambda$, $100\lambda\times 15\lambda$, $150\lambda\times 10\lambda$).
Figure 5: Eigenvalue distribution for different rectangular Tx arrays with the same area: CAP array 1 ($50\lambda\times 30\lambda$), CAP array 2 ($100\lambda\times 15\lambda$), CAP array 3 ($150\lambda\times 10\lambda$). (The broadside linear array $\mathcal{Q}$ has $r_{min}=400\lambda$ and $r_{max}=4000\lambda$.)
1b. Effect of the Rx aperture size:
  • Angular-domain DoF (literature): Similarly to Tx array’s size, DoF scales with the area of Rx array.
  • Distance-domain DoF (this work): Our DoF scales with the inverse-distance span
    $\mathrm{DoF}\;\propto\;\Big(\tfrac{1}{r_{\min}}-\tfrac{1}{r_{\max}}\Big),$
    not directly with the Rx array’s length $r_{\max}-r_{\min}$. As $r_{\max}\to\infty$, the DoF saturates to $\big(p_{\max}^{2}-p_{\min}^{2}\big)\big/(2\lambda r_{\min})$ and does not diverge, unlike typical angular-domain DoF.
1c. Effect of distance:
  • Angular-domain DoF (literature): Element-wise distances between Tx and Rx arrays are nearly equal, so the minimum and maximum Tx–Rx distance are effectively the same, i.e., $r_{\min}\!\approx\!r_{\max}\!\approx\!D$. Hence, the DoF expression typically depends on single value of distance $D$ and $\mathrm{DoF}\sim D^{-n}$.
  • Distance-domain DoF (this work): We allow a large distance spread: it is possible that $r_{\max}\gg r_{\min}$ and $r_{\max}\rightarrow\infty$. The scaling is governed by the two extremes,
    $\mathrm{DoF}\propto\Big(\tfrac{1}{r_{\min}}-\tfrac{1}{r_{\max}}\Big),$
    which captures broad distance spans and saturates to $1/r_{\min}$ as $r_{\max}\to\infty$ (no divergence).
@@ IV-C2 Comparison with Distance-Domain DoF in the Literature
Let us consider [ding2022degrees], which analyzes two perpendicular ULAs—a geometry closest to our setting. If the Tx array center lies on the Rx-line direction, their closed-form DoF can be written as [ding2022degrees, Eqs. (39),(33)]
$\mathrm{DoF}=\frac{D+\tfrac{L_{r}}{2}}{2\lambda}\!\left(\frac{\tfrac{L_{r}}{2}+D}{\sqrt{\big(\tfrac{L_{r}}{2}+D\big)^{2}+\tfrac{L_{t}^{2}}{4}}}-\frac{D-\tfrac{L_{r}}{2}}{\sqrt{\big(D-\tfrac{L_{r}}{2}\big)^{2}+\tfrac{L_{t}^{2}}{4}}}\right),$ (41)
where $D$ is the center-to-center distance between Tx and Rx arrays, and $L_{t},L_{r}$ are the ULA lengths. Using $D=(r_{\max}+r_{\min})/{2},\,L_{r}=r_{\max}-r_{\min},\,p_{\max}=L_{t}/2,$ (41) becomes
$\mathrm{DoF}=\frac{r_{\max}}{2\lambda}\!\left(\frac{r_{\max}}{\sqrt{r_{\max}^{2}+p_{\max}^{2}}}-\frac{r_{\min}}{\sqrt{r_{\min}^{2}+p_{\max}^{2}}}\right).$ (42)
Figure 6: Distance-domain DoF vs length of Rx array obtained via approaches: proposed analytical DoF, [ding2022degrees]. and simulation using SVD (with threshold $\epsilon=0.1\sigma_{\max}$). Tx array is ULA with length of $100\lambda$, and $r_{\min}=100\lambda$, while the length of Rx array ($r_{\max}-r_{\min}$) ranges between $10\lambda,2000\lambda]$.
Narrow distance spans
When $D\gg L_{r}$ (equivalently, $r_{\max}\gg r_{\max}-r_{\min}$), both (42) and our (38) reduce to
$\mathrm{DoF}\;\approx\;\frac{(r_{\max}-r_{\min})\,p_{\max}^{2}}{2\lambda D^{2}},$ (43)
and matches our simulation numerically (see Fig. 6).
Large distance spans
As $r_{\max}\to\infty$, (42) grows without bound, whereas our (38) remains finite and approach:
$\mathrm{DoF}\;\approx\;\frac{\big(p_{\max}^{2}-p_{\min}^{2}\big)}{2\lambda\,r_{\min}},$ (44)
in agreement with simulations (Fig. 6). Thus, when users are spread over a wide range of distances, our proposed DoF expression in (38) captures the correct bounded scaling, while DoF expression in (42) from [ding2022degrees] overestimates growth.
Remark: Unlike many P2P–MIMO DoF’s formulas that implicitly require narrow distance-span assumption ($r_{\min}\!\approx\!r_{\max}$), our distance-domain DoF expression in (38) remains well-posed and can accommodate the large distance spread ($r_{\max}\gg r_{\min}$ or $r_{\max}\rightarrow\infty$). This allows our distance-domain DoF expression to capture the MU-MIMO channel where user’s distance separation can be arbitrarily large.
@@ IV-C3 Non-Uniformity of the Eigenvalue Spectrum
$\widetilde{h}(\boldsymbol{\omega})\;\approx\;\begin{cases}1,&\boldsymbol{\omega}\in\Omega,\\ 0,&\text{otherwise}.\end{cases}$
@@
$\mathrm{DoF}_{\mathrm{eff}}=\frac{(\mathrm{tr}\,\mathbf{V})^{2}}{\mathrm{tr}\,(\mathbf{V}^{\mathsf{H}}\mathbf{V})}=\frac{\big(\sum_{k}\lambda_{k}\big)^{2}}{\sum_{k}\lambda_{k}^{2}},\qquad\mathbf{V}=\mathbf{H}^{\mathsf{H}}\mathbf{H},$ (45)
@@ V-A Transformation to Broadside Case via Projection
Figure 7: Projected Tx array $\bar{\mathcal{P}}$: projection of Tx array $\mathcal{P}$} onto the orthogonal plane to the direction vector $\mathbf{k}$ (For simplicity of illustration, we limit to only x-axis and y-axis).
To simplify the distance-domain DoF analysis of the non-broadside Rx array $\mathcal{Q}$ and Tx array $\mathcal{P}$, we define the projected Tx array $\bar{\mathcal{P}}$, which lies on the plane orthogonal to the direction vector $\mathbf{k}$ of Rx array $\mathcal{Q}$. The projection of a point $\mathbf{p}\in\mathcal{P}$ onto this orthogonal plane is given by:
$\bar{\mathbf{p}}=\left(\mathbf{I}_{3}-\mathbf{k}\mathbf{k}^{T}\right)\mathbf{p},$ (46)
where $\mathbf{I}_{3}$ is the $3\times 3$ identity matrix, $\mathbf{k}\in\mathbb{R}^{3}$ is a unit-norm direction vector. The matrix $\left(\mathbf{I}_{3}-\mathbf{k}\mathbf{k}^{T}\right)$ is defined as the projection matrix that mapped the point into the plane orthogonal to $\mathbf{k}$. Thus, the projected Tx array is defined as:
$\bar{\mathcal{P}}=\left\{\bar{\mathbf{p}}\in\mathbb{R}^{3}\;\middle|\;\bar{\mathbf{p}}=\left(\mathbf{I}_{3}-\mathbf{k}\mathbf{k}^{T}\right)\mathbf{p},\;\mathbf{p}\in\mathcal{P}\right\}.$ (47)
Hence, any arbitrary point $\bar{\mathbf{p}}\in\bar{\mathcal{P}}$ projected from a point $\mathbf{p}\in\mathcal{P}$ satisfies the condition:
$\|\bar{\mathbf{p}}\|^{2}=\|\mathbf{p}\|^{2}-(\mathbf{k}^{T}\mathbf{p})^{2}$ (48)
Subsequently, ${v}(r,r^{\prime})$ in (27) can be rewritten as follows:
@@
We also define the extreme points and extreme distances from the center coordinate $\mathbf{O}$ to $\bar{\mathcal{P}}$ as follows:
$\bar{\mathbf{p}}_{\min}=\operatorname*{argmin}_{\bar{\mathbf{p}}\in\bar{\mathcal{P}}}\|\bar{\mathbf{p}}\|,\quad\bar{\mathbf{p}}_{\max}=\operatorname*{argmax}_{\bar{\mathbf{p}}\in\bar{\mathcal{P}}}\|\bar{\mathbf{p}}\|,$ (50)
By using the projection method, the kernel in (27) of the channel between the Tx array and the Rx array is simplified to the kernel in (49) of the channel between the projected Tx array and the Rx array, where the Rx array lies on the broadside of the Tx array. As a result, by applying Theorem 1 of the broadside case for the projected Tx array and Rx array, we obtain the following DoF:
$\text{DoF}\approx\frac{\left(\bar{p}_{\max}^{2}-\bar{p}_{\min}^{2}\right)}{2\lambda}\left(\frac{1}{r_{\min}}-\frac{1}{r_{\max}}\right)+O(1),$ (51)
Figure 8: Eigenvalue distribution of two cases: 1) Tx array $\mathcal{P}$ and non-broadside Rx array $\mathcal{Q}$ and 2) Projected Tx array $\bar{\mathcal{P}}$ and the same Rx array $\mathcal{Q}$. ($\mathcal{P}$ is rectangular array with the size $150\lambda\times 50\lambda$ and located at the center, while linear array $\mathcal{Q}$ has $\phi=\frac{\pi}{3},\theta=\frac{\pi}{2}$, $r_{\min}=400\lambda$ and $r_{\max}=4000\lambda$).
In the non-broadside configuration, the distance-domain DoF is lower than in the broadside case because the projected Tx array $\bar{\mathcal{P}}$ shrinks as the orientation of $\mathcal{Q}$ deviates from broadside. As illustrated in Fig. 9, when the Rx array $\mathcal{Q}$ is oriented closer to the side of the Tx array $\mathcal{P}$, the projected Tx array $\bar{\mathcal{P}}$ becomes smaller, which in turn reduces the extreme values $\bar{p}_{\max}$ and $\bar{p}_{\min}$ and thus decreases the distance-domain DoF.
Figure 9: Eigenvalue distribution for different angle $\phi$ of Rx array. (The Rx array has $r_{\min}=400\lambda$ and $r_{\max}=4000\lambda$)

### v4

@@ IV-B Eigenvalue Analysis via Fourier Transform Analysis
Scaled Fourier transform defined in (33) of $g(\Delta\zeta)$ in (31) is given by:
@@
Proof: Proof is provided in Appendix B. Note that the Dirac delta function in the definition of $\tilde{g}_{0}(\xi)$ dictates that its non-zero support occurs exclusively where $\xi=-\frac{\|\mathbf{p}\|^{2}}{2\lambda}\big(r^{-1}_{\min}-r^{-1}_{\max}\big)$. Because $r_{\max}>r_{\min}$, the inverse-distance term $(r^{-1}_{\min}-r^{-1}_{\max})$ is strictly positive. Consequently, the spatial-frequency variable $\xi$ is strictly negative.
Figure 3: Geometry of the Tx array $\mathcal{P}$ in the xz-plane (top), the corresponding normalized spectra of $\tilde{g}_{0}(\xi)$ and $\tilde{g}(\xi)$ (middle panels), and eigenvalue distribution (bottom). The squared Tx array spans $60\lambda,100\lambda]$ with a circular gap (radius $60\lambda$). The Rx array spans $200\lambda,2000\lambda]$, and the analytical DoF ignores the $O(1)$ term.
@@
$\text{DoF}=\frac{\left(p_{\max}^{2}-p_{\min}^{2}\right)}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1),$ (38)
where $O(1)$ denotes a bounded residual term that inherently depends on the specific aperture geometry and the chosen eigenvalue threshold.
Note that for scenarios with large distance spans or exceedingly large antenna arrays, the main term grows significantly, making the $O(1)$ term practically negligible. In such regimes, the effective DoF is primarily dominated by the main analytical term, and hence, can be simplified as follows:
$\text{DoF}\approx\left\lceil\frac{\left(p_{\max}^{2}-p_{\min}^{2}\right)}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)\right\rceil,$ (39)
which depends strictly on the aperture’s extreme edges rather than its detailed geometry or the specific eigenvalue threshold.
To explicitly validate this and quantify the $O(1)$ behavior, we simulate the eigenvalue distribution for three distinct aperture shapes sharing the same radial extremes ($p_{\min}=60\lambda,p_{\max}=100\lambda$) and observation span ($r_{\min}=200\lambda,r_{\max}=2000\lambda$). For the feasibility of numerical evaluation, we employ discrete arrays with half-wavelength spacing for both the transmit aperture and the observation array. For this configuration, the main analytical term evaluates to $\left\lceil\frac{\left(p_{\max}^{2}-p_{\min}^{2}\right)}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)\right\rceil=15$. The $O(1)$ term can then be numerically extracted as the residual difference between the simulated effective DoF and this main term. We demonstrate that $O(1)$ fluctuates only slightly across different aperture shapes and thresholds, remaining highly insignificant compared to the main term. For example, applying an effective DoF threshold of $\epsilon=10^{-1}$, the $O(1)$ term evaluates to 0, 11, and $1$ for the rectangular, circular, and linear arrays, respectively. Even when applying a more stringent threshold of $\epsilon=10^{-3}$ to capture a wider transition band, the $O(1)$ term only marginally increases to $1$, 22, and $2$.
It is important to emphasize that despite using continuous arrays as the analytical framework, our proposed analytical DoF accurately predicts the performance of discrete arrays with half-wavelength spacing. To demonstrate this, we numerically evaluated the eigenvalue distributions assuming half-wavelength spacing for different aperture shapes sharing the same extreme edges in Fig. 4, as well as for apertures with different extreme edges in Fig. 5. The results confirm that the numerical DoF obtained from discrete half-wavelength arrays tightly matches our continuous analytical limits.
@@ IV-C1 Comparison with Angular-Domain DoF Expressions
In conventional point-to-point (P2P) MIMO setups, the Rx aperture primarily spans the angular extent of the Tx aperture. Consequently, existing angular-domain DoF expressions typically take the following form [torkildson2011indoor, pizzo2022landau, miller2000communicating, xie2023performance, pu2014effects, ruiz2023degrees, do2023parabolic, xu2017degrees, yuan2021electromagnetic, xu2023exploiting, guo2025impact]:
@@
where $A_{t}$ and $A_{r}$ denote the physical areas of the Tx and Rx arrays, $D$ is the Tx–Rx separation (implicitly treated as a single, uniform distance), $n$ depends on the aperture dimensionality, and $\alpha$ is a coefficient determined by the array orientations. More elaborate variants appear in [wang2025analytical]. Our distance-domain result derived in (38) fundamentally diverges from this standard model in two key aspects:
Figure 4: Three different Tx array geometries in the xz-plane (top panels) and their corresponding normalized eigenvalue distributions in log scale (middle panel) and linear scale (bottom panel). All three shapes share the same radial bounds ($p_{\min}=60\lambda$ and $p_{\max}=100\lambda$). The analytical DoF is obtained via (39).
1) Scaling with Tx Aperture Geometry: Unlike angular-domain DoF, which scales monotonically with the physical array area (e.g., linearly with ULA length or quadratically with URA side length), our distance-domain DoF depends solely on the difference of the squared radial extremes, $\big(p_{\max}^{2}-p_{\min}^{2}\big)$. As demonstrated in Fig. 4, arrays with vastly different detailed shapes and physical areas yield nearly identical distance-domain DoF provided they share the same radial boundaries $(p_{\max},p_{\min})$. Consequently, for a fixed aperture area, strategically elongating the array to maximize $p_{\max}$ is far more effective for increasing the DoF than uniformly enlarging the area. This geometric scaling advantage is explicitly verified in Fig. 5, which compares three rectangular arrays with identical physical areas but varying elongations.
Figure 5: Geometries of three rectangular Tx arrays sharing the exact same physical area (top panels) and their corresponding normalized eigenvalue distributions (bottom panel). The array dimensions are $50\lambda\times 30\lambda$ (Array 1), $100\lambda\times 15\lambda$ (Array 2), and $150\lambda\times 10\lambda$ (Array 3). The broadside linear Rx array $\mathcal{Q}$ spans from $r_{\min}=400\lambda$ to $r_{\max}=4000\lambda$.
2) Scaling with Rx Span and Distance Spread: Angular-domain models implicitly assume a narrow spatial span ($r_{\min}\approx r_{\max}\approx D$), causing the DoF to scale simply with the receive area and $D^{-n}$. In stark contrast, our framework accurately models systems with exceptionally large distance spreads ($r_{\max}\gg r_{\min}$). The distance-domain DoF scales with the inverse-distance span, $\big(\frac{1}{r_{\min}}-\frac{1}{r_{\max}}\big)$, rather than the absolute physical length of the Rx array. A profound physical consequence of this scaling is that as the receive observation region extends toward infinity ($r_{\max}\to\infty$), the DoF does not exhibit unphysical, unbounded growth; instead, it safely saturates to a finite limit bounded by $1/r_{\min}$.
@@ IV-C2 Comparison with Distance-Domain DoF in the Literature
We compare our framework with [ding2022degrees], which analyzes two perpendicular ULAs—the geometry closest to our broadside setting. By mapping their geometric variables (center-to-center distance $D$, and ULA lengths $L_{t},L_{r}$) to our radial notation ($D=(r_{\max}+r_{\min})/{2}$, $L_{r}=r_{\max}-r_{\min}$, and $p_{\max}=L_{t}/2$), their closed-form DoF expression [ding2022degrees, Eqs. (39),(33)] translates directly to:
[…]
Figure 6: Distance-domain DoF vs. length of Rx array obtained via: the proposed analytical DoF, the expression from [ding2022degrees], and SVD simulations (threshold $\epsilon=0.1\sigma_{\max}$). The Tx array is a ULA of length $100\lambda$, $r_{\min}=100\lambda$, and the Rx array length ($r_{\max}-r_{\min}$) ranges between $10\lambda,2000\lambda]$.
As illustrated in Fig. 6, the validity of (42) depends heavily on the distance span. For narrow distance spans ($D\gg L_{r}$, or $r_{\max}\approx r_{\min}$), both (42) and our proposed expression in (38) reduce to the same approximation, $\mathrm{DoF}\approx\frac{(r_{\max}-r_{\min})p_{\max}^{2}}{2\lambda D^{2}}$, tightly matching the numerical simulations.
However, a critical divergence occurs for large distance spans. As $r_{\max}\to\infty$, the expression from [ding2022degrees] grows without bound. In stark contrast, our proposed DoF remains physically finite, correctly saturating to $\frac{p_{\max}^{2}-p_{\min}^{2}}{2\lambda\,r_{\min}}$. Thus, when users or observation regions are spread over a wide range of distances, our framework captures the correct bounded scaling, whereas the prior model in [ding2022degrees] severely overestimates the available DoF.
Remark: To summarize, our framework overcomes the fundamental limitations of both existing angular-domain and distance-domain DoF models in [ding2022degrees]. While conventional angular-domain models scale heavily with aperture area, and prior distance-domain works like [ding2022degrees] are restricted to specific array geometries and narrow distance spans ($r_{\min}\!\approx\!r_{\max}$), our derived expression in (38) offers a universal solution. It applies to arbitrary two-dimensional transmit apertures by depending solely on radial extremes, and it remains strictly well-posed for unbounded distance spreads ($r_{\max}\rightarrow\infty$). Consequently, our model uniquely captures the spatial multiplexing capabilities of realistic near-field MU-MIMO networks where both the base station aperture shape and user distance separations can be highly generalized.
@@ IV-C3 Non-Uniformity of the Eigenvalue Spectrum
$\widetilde{h}(\bm{\omega})\;\approx\;\begin{cases}1,&\bm{\omega}\in\Omega,\\ 0,&\text{otherwise}.\end{cases}$
@@
$\text{DoF}_{\text{eff}}=\frac{(\text{tr}\,\mathbf{V})^{2}}{\text{tr}\,(\mathbf{V}^{\text{H}}\mathbf{V})}=\frac{\big(\sum_{k}\lambda_{k}\big)^{2}}{\sum_{k}\lambda_{k}^{2}},\qquad\mathbf{V}=\mathbf{H}^{\text{H}}\mathbf{H},$ (43)
@@ V-A Transformation to Broadside Case via Projection
Figure 7: Projected Tx array $\bar{\mathcal{P}}$: projection of Tx array $\mathcal{P}$} onto the plane orthogonal to the direction vector $\mathbf{k}$ (For simplicity of illustration, we limit to only the x- and y-axes).
To analyze the non-broadside configuration efficiently, we project the Tx array $\mathcal{P}$ onto the plane orthogonal to the Rx array’s direction vector $\mathbf{k}$. Using the projection matrix $\mathbf{I}_{3}-\mathbf{k}\mathbf{k}^{T}$, the projected Tx array is defined concisely as:
$\bar{\mathcal{P}}=\left\{\bar{\mathbf{p}}=\left(\mathbf{I}_{3}-\mathbf{k}\mathbf{k}^{T}\right)\mathbf{p}\;\middle|\;\mathbf{p}\in\mathcal{P}\right\},$ (44)
where any projected point $\bar{\mathbf{p}}\in\bar{\mathcal{P}}$ satisfies the geometric condition $\|\bar{\mathbf{p}}\|^{2}=\|\mathbf{p}\|^{2}-(\mathbf{k}^{T}\mathbf{p})^{2}$. Substituting this relationship into the non-broadside channel model, the corresponding kernel function can be rewritten as an integral over the projected aperture $\bar{\mathcal{P}}$:
@@
Crucially, this kernel function is exactly similar to the broadside case derived in (28), with the original coordinates $\mathbf{p}$ and aperture $\mathcal{P}$ simply replaced by their projected counterparts $\bar{\mathbf{p}}$ and $\bar{\mathcal{P}}$.
Because the mathematical structure is identical, we can directly apply Theorem 1 to this equivalent broadside system. The distance-domain DoF is obtained by simply substituting the original radial boundaries in (38) with the extreme radial distances of the projected array ($\bar{p}_{\min}$ and $\bar{p}_{\max}$):
$\text{DoF}\approx\frac{\left(\bar{p}_{\max}^{2}-\bar{p}_{\min}^{2}\right)}{2\lambda}\left(\frac{1}{r_{\min}}-\frac{1}{r_{\max}}\right)+O(1).$ (46)
Figure 8: Eigenvalue distribution comparison: 1) Tx array $\mathcal{P}$ and non-broadside Rx array $\mathcal{Q}$, versus 2) Projected Tx array $\bar{\mathcal{P}}$ and the same Rx array $\mathcal{Q}$. ($\mathcal{P}$ is a $150\lambda\times 50\lambda$ rectangular array located at the origin; the linear array $\mathcal{Q}$ has $\phi=\frac{\pi}{3},\theta=\frac{\pi}{2}$, with $r_{\min}=400\lambda$ and $r_{\max}=4000\lambda$).
As verified by the eigenvalue distributions in Fig. 8, this projected broadside model perfectly characterizes the true non-broadside channel. Physically, the distance-domain DoF in the non-broadside configuration is fundamentally lower than in the broadside case. As illustrated in Fig. 9, when the Rx array $\mathcal{Q}$ is oriented further away from the broadside direction, the effective projected Tx array $\bar{\mathcal{P}}$ shrinks. This geometric reduction directly decreases the extreme radial values $\bar{p}_{\max}$ and $\bar{p}_{\min}$, which in turn restricts the available distance-domain DoF.
Figure 9: Original Tx array aperture (top panel) and its projections for different Rx array angles $\phi$ (middle panels) and the corresponding eigenvalue distributions (bottom panel) of the corresponding channels. The Rx array spans from $r_{\min}=400\lambda$ to $r_{\max}=4000\lambda$.